convert But-2-ene  to 2-methylbut-2-enol.kindly answer it quickly

Grade:11

1 Answers

ALI JINNAH
8 years ago
H3C-CH=CH-CH3--------> H3C-CH=C-CH2-OH
                                                          !
                                                          CH3
In this eqution 2-butene is converted into 2-methylbut-2-enol.